# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ 4.2. Improved LWP for mZK
and due to $\lvert\xi_{1}+\xi_{3}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, $\lvert\xi_{2}+\xi_{3}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, (49), and the inequalities in (50), we obtain in the case $\lvert\xi_{1}+\xi_{2}\rvert\gtrsim 1$ that
$\lvert R\rvert\gtrsim\lvert\xi_{\text{max}}\rvert^{2},$
which, as in case (ii.3.2.2), is already sufficient to obtain (42) for every $s>\frac{1}{3}$. Conversely, in the case $\lvert\xi_{1}+\xi_{2}\rvert\ll 1$, it follows from (50) that $\lvert q_{1}+q_{2}\rvert\ll 1$ as well. We now define, for dyadic $N\in 2^{\mathbb{N}_{0}}$ and $\beta\in\mathbb{R}$, the Fourier projectors $Q_{N}^{(\beta)}$ by setting
Taking into account the estimate
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\xi_{0}\rvert^{\frac{1}{8}}\prod_{i=1}^{3}\lvert\xi_{i}\rvert^{\frac{1}{8}}\langle(\xi_{i},q_{i})\rangle^{\frac{s}{3}+\frac{1}{6}},$
the dual version of (19), followed by Hölder’s inequality then yields
$\lVert\partial_{x}(Q_{N}^{(0)}(u_{1}u_{2})u_{3})\rVert_{X_{s,-\frac{1}{2}+2\varepsilon}}$
$\lesssim\lVert I_{x}^{\frac{1}{8}}(Q_{N}^{(0)}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{3})\rVert_{X_{0,-\frac{1}{2}+2\varepsilon}}$
$\lesssim\lVert Q_{N}^{(0)}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{3}\rVert_{L_{tx}^{\frac{4}{3}}L_{y}^{2}}$
$\leq\lVert Q_{N}^{(0)}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})\rVert_{L_{tx}^{2}L_{y}^{\infty}}\lVert I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{3}\rVert_{L_{tx}^{4}L_{y}^{2}},$
$\lVert Q_{N}^{(0)}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})\rVert_{L_{y}^{\infty}}$ $\lesssim\lVert\chi_{\{\lvert q\rvert\lesssim 1\}}\mathcal{F}_{y}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})\rVert_{L_{q}^{1}}$
$\leq\lVert\chi_{\{\lvert q\rvert\lesssim 1\}}\rVert_{L_{q}^{1}}\lVert\mathcal{F}_{y}(I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2})\rVert_{L_{q}^{\infty}}$
$\lesssim\lVert I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2}\rVert_{L_{y}^{1}}.$
With this, we further obtain
$\lVert\partial_{x}(Q_{N}^{(0)}(u_{1}u_{2})u_{3})\rVert_{X_{s,-\frac{1}{2}+2\varepsilon}}$ $\lesssim\lVert I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{1}I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{2}\rVert_{L_{tx}^{2}L_{y}^{1}}\lVert I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{3}\rVert_{L_{tx}^{4}L_{y}^{2}}$
$\leq\prod_{i=1}^{3}\lVert I_{x}^{\frac{1}{8}}J^{\frac{s}{3}+\frac{1}{6}}u_{i}\rVert_{L_{tx}^{4}L_{y}^{2}},$
and three applications of (19) ultimately give
... $\lesssim\prod_{i=1}^{3}\lVert u_{i}\rVert_{X_{\frac{s}{3}+\frac{1}{6},\frac{1}{2}+\varepsilon}}$
$\lesssim\prod_{i=1}^{3}\lVert u_{i}\rVert_{X_{s,\frac{1}{2}+\varepsilon}},$
where the final step requires
$\frac{s}{3}+\frac{1}{6}\leq s\Leftrightarrow\frac{1}{4}\leq s.$
For sufficiently small $\varepsilon>0$, this completes the proof of (42) for all $s\geq\frac{1}{4}$ in this situation.
(ii.3.3.3) $\lvert\xi_{i}+\xi_{j}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, for exactly one $(i,j)\in\{(1,2),(1,3),(2,3)\}$:
Without loss of generality, we may assume that $\lvert\xi_{1}+\xi_{2}\rvert\ll\lvert\xi_{\text{max}}\rvert$ and $\lvert\xi_{2}+\xi_{3}\rvert\ll\lvert\xi_{\text{max}}\rvert$. Then $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{1}+\xi_{3}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{1}\rvert$, and $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{3}\rvert$ hold, and if the bilinear refinement is applicable in one of these situations, we can argue exactly as in case (ii.3.1) and obtain the desired trilinear estimate for every $s>\frac{3}{8}$. Therefore, we have to consider the situation where
$\begin{split}\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert&\ll\lvert\xi_{1}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{1}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert\end{split}$ (52)
hold simultaneously. From the assumptions (ii.3), (ii.3.3), $\lvert\xi_{1}+\xi_{2}\rvert\ll\lvert\xi_{\text{max}}\rvert$,
$\lvert\xi_{2}+\xi_{3}\rvert\ll\lvert\xi_{\text{max}}\rvert$, and (52) it can now be easily deduced that
$\lvert\xi_{0}\rvert\sim\lvert\xi_{1}\rvert\sim\lvert\xi_{2}\rvert\sim\lvert\xi_{3}\rvert\sim\lvert q_{0}\rvert\sim\lvert q_{1}\rvert\sim\lvert q_{2}\rvert\sim\lvert q_{3}\rvert\sim\lvert(\xi_{1},q_{1})\rvert$ (53)
@@
$\lvert\lvert\xi_{i}\rvert-\lvert\xi_{j}\rvert\rvert\ll\lvert\xi_{0}\rvert\quad\text{and}\quad\lvert\lvert q_{i}\rvert-\lvert q_{j}\rvert\rvert\ll\lvert q_{0}\rvert,\quad\forall i,j\in\{0,1,2,3\},$ (54)
and these relations impose the following possible sign constellations.
$(\xi_{0},\xi_{1},\xi_{2},\xi_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
In what follows, we may once again restrict our attention to a specific case, say $(\xi_{0},\xi_{1},\xi_{2},\xi_{3})=(+,+,-,+)$ and $(q_{0},q_{1},q_{2},q_{3})=(-,-,+,-)$, as the remaining cases proceed analogously. Substituting this sign configuration into (52) and taking (53) into account, we arrive at
$\lvert\sqrt{3}(\xi_{1}+\xi_{3})+(q_{1}+q_{3})\rvert\ll 1,$
$\lvert\sqrt{3}(\xi_{0}+\xi_{1})+(q_{0}+q_{1})\rvert\ll 1,$
and
$\lvert\sqrt{3}(\xi_{0}+\xi_{3})+(q_{0}+q_{3})\rvert\ll 1,$
and adding and subtracting these inequalities leads to
$\begin{split}\lvert\sqrt{3}\xi_{0}+q_{0}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{1}+q_{1}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{2}+q_{2}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{3}+q_{3}\rvert&\ll 1.\end{split}$ (55)
In contrast to case (ii.3.3.2), these estimates are no longer sufficient to ensure that the resonance function of mKdV dominates the remaining terms in (51). To enforce this, we now consider the following alternatives:
$\begin{split}(\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert\gtrsim\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{2}}\ &\text{and}\ \lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{5}{4}})\\ \text{or}\ (\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{1}\rvert^{\frac{1}{2}}\ &\text{and}\ \lvert\lvert(\xi_{2},q_{2})\rvert^{2}-\lvert(\xi_{3},q_{3})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{5}{4}})\\ \text{or}\ (\lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{3}\rvert^{\frac{1}{2}}\ &\text{and}\ \lvert\lvert(\xi_{1},q_{1})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{5}{4}}).\end{split}$
We show that the first two assumptions each lead to the desired result. The argument for the third assumption is the same as for the second.
If $\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert\gtrsim\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{2}}$ and $\lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{5}{4}}$, then we have
@@
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{8}}\langle(\xi_{1},q_{1})\rangle^{\frac{s}{2}+\frac{3}{16}+}\langle(\xi_{3},q_{3})\rangle^{\frac{s}{2}+\frac{3}{16}+}$
$\ \ \ \cdot\lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert^{\frac{1}{2}}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{2},q_{2})\rangle^{-\frac{1}{8}},$
and undoing Plancherel, followed by Hölder’s inequality yields
$I_{f}\lesssim\lVert I_{x}^{\frac{1}{8}}P^{\frac{1}{2}}(J^{\frac{s}{2}+\frac{3}{16}+}u_{1}J^{\frac{s}{2}+\frac{3}{16}+}u_{3})\rVert_{L_{txy}^{2}}\lVert MP(J^{0-}f,J^{-\frac{1}{8}}\tilde{u}_{2})\rVert_{L_{txy}^{2}}.$
For the first factor, we can now invoke the bilinear refinement (30), and for the second factor, we apply the bilinear estimate (5). It then follows that
... $\lesssim\lVert u_{1}\rVert_{X_{\frac{s}{2}+\frac{3}{16}+,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{s}{2}+\frac{3}{16}+,\frac{1}{2}+\varepsilon}}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}$
$\lesssim\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}},$
and the last step requires
$\frac{s}{2}+\frac{3}{16}+\leq s\Leftrightarrow\frac{3}{8}+\leq s,$
which, with a suitable choice of $0<\varepsilon\ll 1$, establishes the desired trilinear estimate (42) for all $s>\frac{3}{8}$. On the other hand, under the assumptions
$\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{1}\rvert^{\frac{1}{2}}$ and $\lvert\lvert(\xi_{2},q_{2})\rvert^{2}-\lvert(\xi_{3},q_{3})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{5}{4}}$, we have
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\xi_{0}+\xi_{1}\rvert^{\frac{1}{8}}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{1},q_{1})\rangle^{\frac{3}{8}+}$
$\ \ \ \cdot\lvert\lvert(\xi_{2},q_{2})\rvert^{2}-\lvert(\xi_{3},q_{3})\rvert^{2}\rvert^{\frac{1}{2}}\langle(\xi_{2},q_{2})\rangle^{s}\langle(\xi_{3},q_{3})\rangle^{-\frac{1}{8}},$
and this time undoing Plancherel, followed by Hölder’s inequality yields
$I_{f}\lesssim\lVert I_{x}^{\frac{1}{8}}P^{\frac{1}{2}}(J^{0-}fJ^{\frac{3}{8}+}\tilde{u}_{1})\rVert_{L_{txy}^{2}}\lVert MP(J^{s}u_{2},J^{-\frac{1}{8}}u_{3})\rVert_{L_{txy}^{2}}.$
The first factor is now handled, using (33), while for the second factor we can invoke (3). We thus obtain
... $\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{1}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{2}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}},$
so that choosing $\varepsilon>0$ sufficiently small completes the discussion of this case as well. It remains to consider what happens in the case
$\begin{split}(\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert\ll\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{2}}\ &\text{or}\ \lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\ll\lvert\xi_{0}\rvert^{\frac{5}{4}})\\ \text{and}\ (\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert\ll\lvert\xi_{0}+\xi_{1}\rvert^{\frac{1}{2}}\ &\text{or}\ \lvert\lvert(\xi_{2},q_{2})\rvert^{2}-\lvert(\xi_{3},q_{3})\rvert^{2}\rvert\ll\lvert\xi_{0}\rvert^{\frac{5}{4}})\\ \text{and}\ (\lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert\ll\lvert\xi_{0}+\xi_{3}\rvert^{\frac{1}{2}}\ &\text{or}\ \lvert\lvert(\xi_{1},q_{1})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\ll\lvert\xi_{0}\rvert^{\frac{5}{4}}).\end{split}$
If $\lvert\lvert(\xi_{i},q_{i})\rvert^{2}-\lvert(\xi_{j},q_{j})\rvert^{2}\rvert\ll\lvert\xi_{0}\rvert^{\frac{5}{4}}$ holds for some pair $(i,j)\in\{(0,2),(2,3),(1,2)\}$, then in view of (55), we obtain
$\lvert 2(q_{i}^{2}-q_{j}^{2})\rvert$ $=\lvert 3\xi_{i}^{2}+q_{i}^{2}-(3\xi_{j}^{2}+q_{j}^{2})+q_{i}^{2}-3\xi_{i}^{2}-q_{j}^{2}+3\xi_{j}^{2}\rvert$
$\leq\lvert\lvert(\xi_{i},q_{i})\rvert^{2}-\lvert(\xi_{j},q_{j})\rvert^{2}\rvert+\lvert\sqrt{3}\xi_{i}+q_{i}\rvert\lvert\sqrt{3}\xi_{i}-q_{i}\rvert$
$\ \ \ +\lvert\sqrt{3}\xi_{j}+q_{j}\rvert\lvert\sqrt{3}\xi_{j}-q_{j}\rvert$
$\ll\lvert\xi_{0}\rvert^{\frac{5}{4}};$
and since $q_{i}$ and $q_{j}$ have opposite signs for the indices under consideration, it follows that
$\lvert q_{i}+q_{j}\rvert\ll\lvert\xi_{0}\rvert^{\frac{1}{4}}.$
Now, setting
$f=P_{N}f,\ u_{1}=P_{N_{1}}u_{1},\ u_{2}=P_{N_{2}}u_{2},\ \text{and}\ u_{3}=P_{N_{3}}u_{3}$
for dyadic numbers $N,N_{1},N_{2},N_{3}\in 2^{\mathbb{N}_{0}}$, it follows that $\lvert q_{i}+q_{j}\rvert\ll N^{\frac{1}{4}}$ and by redistributing derivatives, followed by an application of Hölder’s inequality, we obtain
$I_{f}\lesssim\sum_{\begin{subarray}{c}N\sim N_{1}\sim N_{2}\sim N_{3}\\ N\gg 1\end{subarray}}N^{\frac{1}{2}-2s}\lVert Q_{N}^{(\frac{1}{4})}(I_{x}^{\frac{1}{8}}fI_{x}^{\frac{1}{8}}J^{s}\tilde{u}_{2})I_{x}^{\frac{1}{8}}J^{s}\tilde{u}_{1}\rVert_{L_{tx}^{\frac{4}{3}}L_{y}^{2}}\lVert I_{x}^{\frac{1}{8}}J^{s}u_{3}\rVert_{L_{tx}^{4}L_{y}^{2}}$
in the case $(i,j)=(0,2)$, and

### v3

@@ 4.2. Improved LWP for mZK
In this case, however, since ${\lvert}\xi_{1}+\xi_{2}{\rvert},{\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}$, we obtain
${\lvert}\xi_{1}+\xi_{2}{\rvert}{\lvert}\xi_{2}+\xi_{3}{\rvert}{\lvert}\xi_{1}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{13}{6}},$
and together with (ii.3) this yields
${\lvert}R{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{13}{6}}.$
As previously discussed, this already suffices to obtain at least $s>\frac{1}{3}$, so that we can also assume ${\lvert}q_{1}-q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}$. Moreover, these considerations work entirely analogously for the pairing $q_{0}+q_{2}$, which allows us to include ${\lvert}q_{0}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}$ among our remaining active constraints as well. As a final preparatory step for our last argument, we write
$f=P_{N}f,\ u_{1}=P_{N_{1}}u_{1},\ u_{2}=P_{N_{2}}u_{2},\ \text{and}\ u_{3}=P_{N_{3}}u_{3}$
for dyadic numbers $N\sim N_{1}\sim N_{2}\sim N_{3}\gg 1$, which allows us to summarize all the relevant remaining constraints in the form:
${\lvert}\sqrt{3}(\xi_{1}+\xi_{3})\mp(q_{1}+q_{3}){\rvert}\ll N^{-\frac{5}{6}},\ {\lvert}q_{1}-q_{3}{\rvert}\ll N^{\frac{3}{4}},\ \text{and}\ {\lvert}q_{0}+q_{2}{\rvert}\ll N^{\frac{3}{4}}.$ (45)
Let us now define, for $\alpha,\beta\in\mathbb{R}$ and dyadic $N\gg 1$, the bilinear restriction operator $S_{N}^{(\alpha,\beta)}$ by means of its Fourier transform:
$\widehat{S_{N}^{(\alpha,\beta)}(u,v)}(\tau,\xi,q)\coloneqq$
$\int_{\mathbb{R}^{2}}\sum_{q_{1}\in\mathbb{Z}}\chi_{\{{\lvert}\sqrt{3}\xi\mp q{\rvert}\lesssim N^{\alpha}\}}\chi_{\{{\lvert}2q_{1}-q{\rvert}\lesssim N^{\beta}\}}\hat{u}(\tau_{1},\xi_{1},q_{1})\hat{v}(\tau-\tau_{1},\xi-\xi_{1},q-q_{1})\mathrm{d}(\tau_{1},\xi_{1}).$
Then, taking into account
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}\lesssim N^{1-2s+}\langle(\xi_{1},q_{1})\rangle^{s}\langle(\xi_{3},q_{3})\rangle^{s}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{2},q_{2})\rangle^{s}$
and the constraints in (45), it follows - by undoing Plancherel and an application of Hölder’s inequality - that
$I_{f}\lesssim N^{1-2s+}{\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(J^{s}u_{1},J^{s}u_{3}){\rVert}_{L_{txy}^{2}}{\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(J^{0-}f,J^{s}\tilde{u}_{2}){\rVert}_{L_{txy}^{2}},$
and it remains to appropriately estimate the two $L^{2}$-norms. To this end, let $b>\frac{1}{2}$ be given arbitrarily. Then, by duality, and after applying Plancherel’s theorem, followed by the Cauchy-Schwarz inequality and Fubini’s theorem, we obtain
${\lVert}S_{N}^{(-\frac{5}{6},\frac{3}{4})}(u,v){\rVert}_{L_{txy}^{2}}\lesssim\Bigg(\sup_{(\tau_{1},\xi_{1},q_{1})\in\mathbb{R}\times\mathbb{R}\times\mathbb{Z}}c_{0}^{\frac{1}{2}}\Bigg){\lVert}u{\rVert}_{X_{0,b}}{\lVert}v{\rVert}_{X_{0,b}}$
@@
$c_{0}$ $=\langle\tau_{1}-\varphi(\xi_{1},q_{1})\rangle^{-2b}$
@@
$\ \ \ \cdot\sum_{\begin{subarray}{c}q\in\mathbb{Z}\\ {\lvert}2q_{1}-q{\rvert}\lesssim N^{\frac{3}{4}}\end{subarray}}\int_{\mathbb{R}}\chi_{\{{\lvert}\sqrt{3}\xi\mp q{\rvert}\lesssim N^{-\frac{5}{6}}\}}\int_{\mathbb{R}}\langle\tau-\tau_{1}-\varphi(\xi-\xi_{1},q-q_{1})\rangle^{-2b}\mathrm{d}\tau\mathrm{d}\xi,$
and since $b>\frac{1}{2}$, the integral over $\tau$ is finite. Consequently, we have